# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v1

@@ 4. Parabolic restriction and difference operators
In this section, we recall the notion of the parabolic restriction for Harish-Chandra bimodules essentially following [KalmykovSafronov], as well as the classical construction of Mickelsson algebras [ZhelobenkoSalgebras]. In particular, we identify an example of the latter with a negative half of the algebra of difference operators.
@@ 4.2. Parabolic restriction
In this subsection, we recall the notion of parabolic restriction as in [KalmykovSafronov] and explicitly compute in some concrete examples.
@@
In fact, one can explicitly compute the parabolic restriction as follows. By [AsherovaSmirnovTolstoy], there exists an element $P$ of $\mathfrak{h}_{N}$-weight zero in a certain completion of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}[\hbar^{-1}]$ whose action is well-defined on any right $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}$-module $F$ with a locally-finite action of $\mathfrak{n}_{-}$ satisfying $F\mathfrak{n}_{-}\subset\hbar F$; this element is called the extremal projector. It satisfies the following properties:
[…]
@@
Usually, the cited results use the positive version of the parabolic restriction $X\mapsto(X^{\mathrm{gen}}/\mathfrak{n}_{+})^{N_{+}}$. For Miura transform of Section 6, the negative one will be more convenient as the former is a map of left $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-modules; however, both versions are equivalent thanks to the property $\varpi_{N}(P)=P$. For instance, assume $\varpi_{N}$ can be extended to an isomorphism of a Harish-Chandra bimodule $X$. Then there is an anti-isomorphism
@@
where, as usual, the first part is the image of $A(u)$ in the quotient $\mathfrak{n}_{-}\backslash\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$ and the second part is its image in the quotient $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\mathfrak{n}_{+}$ by the right action of $\mathfrak{n}_{+}$, see [MolevYangians, Theorem 7.1.1]. In particular, this defines an isomorphism $\mathbf{Z}_{\mathfrak{gl}_{N}}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{W}$, where $W=S_{N}$ acts by permutations of coordinates $\{w_{i}\}$. In terms of coordinates $\{E_{ii}\}$, it corresponds to the $\rho_{N}$-shifted action
@@
This is the standard Harish-Chandra isomorphism $\mathbf{Z}_{\mathfrak{gl}_{N}}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{W}$, see [HumphreysBGG, Theorem 1.10].
@@ 4.3. Vector representation
where $\{j_{1}\ldots j_{r}\}$ is the complementary subset to $\{i_{1},\ldots,i_{s}\}$ inside $\{i+1,\ldots,N\}$. According to formulas [MolevTransvectorAlgebras, (3.2),(3.3)], the trick of Example 3.6, and Remark 4.3, those vectors are indeed invariant under $N_{-}$ and satisfy
@@
Using the parabolic restriction, one can deduce the following explicit description of the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$ as bimodule over the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$. Before the statement, we need the following lemma which is essentially a part of [KhoroshkinNazarovVinberg, Theorem 2]; for reader’s convenience, we repeat its proof in our notations.
@@
Recall that parabolic restriction is isomorphic to the functor $X\mapsto\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}_{+}$ as in (4.7). For the category $\mathrm{QCoh}_{[\mathfrak{g}^{*}/G]}$ as a subcategory of Harish-Chandra bimodules on which $\hbar$ acts by zero as in Remark 3.3, this functor is simply the restriction along the embedding $(\mathfrak{h}_{N}^{*})^{\mathrm{reg}}\hookrightarrow\mathfrak{g}^{*}$. Since its $\mathrm{GL}_{N}$-orbit is dense, the restriction $\mathrm{res}(s)$ is non-zero as well. Since the parabolic restriction functor commutes with $\mathbf{C}[\hbar]$-action, we obtain that $\mathrm{res}(s_{\hbar})$ is non-zero too. ∎
@@
The equality in the following proposition is understood as in Remark 4.8.
@@
Observe that both sides are actions of certain elements of $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})$. By 4.5, it is enough to show the equality under the parabolic restriction functor. It follows from 4.6 that
@@
At the same time, recall that by Example 4.4, we have $PA(u)=\prod_{j=1}^{N}(u-w_{j})$, therefore,
@@
Strictly speaking, we understand the equality of Proposition 4.7 in the sense of power series expansion in $u^{-1}$. The right-hand side can be presented as
@@
while the left-hand side can be equivalently presented as conjugation by $\mathrm{qdet}(T(-u+N\hbar-\hbar))$ in the notations of Section 1.2 which also defines an element of $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})[[u^{-1}]]$. Another way is via the coefficients (1.11): for any $v\in\mathbf{C}^{N}$, we have
@@
In fact, using 4.6 and (4.9), one can re-deduce the Cayley-Hamilton theorem of Proposition 3.7: observe that for every $i$, we have
@@ 4.4. Mickelsson algebras and difference operators
By [KalmykovSafronov, Corollary 4.18], the natural lax monoidal structure on $\mathrm{res}$ given by tensor multiplication
@@
[ZhelobenkoSalgebras] Let $A\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$ be an algebra object. The Mickelsson algebra associated to $A$ is the $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodule $A^{\mathrm{gen}}/\!\!/N_{-}$ with the multiplication given by
@@
where the target is the space of polynomials of total degree $k$ in $\mathbf{u}^{-1}$ variables. By 3.9, the source has an action of the spherical subalgebra $e^{\hbar}_{k}\mathfrak{H}^{\hbar}_{k}e^{\hbar}_{k}$ of the degenerate affine Hecke algebra.
@@
Recall that $\mathbf{u}_{i_{1}}^{-1}\ldots\mathbf{u}_{i_{k}}^{-1}$ is proportional to $Pv_{i_{1}}P\ldots Pv_{i_{k}}P$. By 3.9, the action of a monomial on the latter is
@@
(it is well-defined once the function is symmetric). The result follows by 4.6. ∎
@@ 5. Kostant-Whittaker reduction and Yangians
In this section, we will study a variant of the constructions of Section 4. Our main goal is to relate the shifted Yangian of $\mathfrak{sl}_{2}$ to the Whittaker analog of the Mickelsson algebra for $S^{\bullet}\mathbf{C}^{N}$.
@@
For any $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$, denote by $X/\mathfrak{n}_{+}^{\psi}$ the quotient by the right action of $\mathfrak{n}_{+}^{\psi}$. Recall that by Kostant’s construction [Kostant], the space of invariants $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\!\!/_{\psi}\,N_{+}:=(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\mathfrak{n}_{+}^{\psi})^{N_{+}}$ is isomorphic to the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$. It admits the following categorification.
@@
[BezrukavnikovFinkelberg] The Kostant-Whittaker reduction is a functor
@@
There is a Whittaker analog of the extremal projector (4.6) introduced in [KalmykovYangians] under the name of the Kirillov projector.
@@
[…]
@@
In the notations of [KalmykovYangians, Definition 5.2], we can apply (1.6) to $P_{\mathfrak{m}_{N}}^{\psi}(N\hbar-\hbar,\ldots,\hbar)$; by [KalmykovYangians, Theorem 5.3], it satisfies the properties of the theorem for $\psi(E_{ij})=\delta_{i+1,j}$. For general $\psi\in\mathbf{C}^{*}$, we can apply the automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$ defined by $x\mapsto x/\psi$ for $x\in\mathfrak{gl}_{N}$ and $\hbar\mapsto\hbar/\psi$. ∎
As in Section 4, we use the following.
@@
We can compute explicitly this trivialization for the vector representation $\mathbf{C}^{N}$ using the canonical homomorphism (3.3). The following can be proved analogously to [KalmykovYangians, Proposition 6.18].
@@
Combining (5.3), 5.4 and Proposition 3.7, we conclude that the map
@@
is an isomorphism of modules over $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]$. By Proposition 4.7, we have
@@
By expanding in powers of $u$ as in Remark 4.8, we can conclude that an endomorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ commuting with the left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-action commutes with the right $\mathbf{Z}_{\mathfrak{gl}_{N}}$-action if and only if it commutes with $\Omega$. Therefore, such an endomorphism commutes with the algebra action of $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))$, giving a canonical isomorphism
@@ 5.1. Kostant-Mickelsson-Whittaker algebras
see [BezrukavnikovFinkelberg]. In particular, we have an analog of 4.10.
@@
Recall the action of the degenerate affine Hecke algebra $\mathfrak{H}^{\hbar}_{k}$ on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ from Section 3.1. It descends to an action on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}$. Recall the central elements $A(u)=\sum_{i}A_{i}u^{N-i}$ of (1.11). Define the elements $A^{(j)}_{i}\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]$ for $11\leq j\leq k+$ 1 inductively via $A^{(1)}_{i}=A_{i}$ for all $i$ and
@@
Using Remark 4.8, it can be rewritten as
@@
The following is a generalization of Proposition 5.5.
@@
By Proposition 4.7, the right action of the center on each tensor component satisfies
@@
By A.1, they generate the ring of symmetric functions in $\{\Omega_{i}\}$. Therefore, we have a desired isomorphism. ∎
@@
The following theorem relates the positive part of the shifted Yangian from 2.13 to the Kostant-Mickelsson-Whittaker algebra associated to $S^{\bullet}\mathbf{C}^{N}$. Observe similarity of the explicit formulas below to the quantization of Zastava spaces for $\mathrm{SL}_{2}$ constructed in [FinkelbergRybnikov].
@@
$-A(z)\frac{1}{w-\Omega_{1}}(v_{1})+\frac{1}{w-\Omega_{1}}(v_{1})A(z)=-\frac{\hbar}{z-w}A(z)\left(\frac{1}{z-\Omega_{1}}-\frac{1}{w-\Omega_{1}}\right)(v_{1}).$
@@
and, due to Proposition 4.7, the left-hand side is
$-A(z)\frac{1}{w-\Omega_{1}}(v_{1})+A(z)\frac{1}{w-\Omega_{1}}(v_{1})\left(1+\frac{\hbar}{z-\Omega_{1}}\right)=A(z)\frac{\hbar}{(z-\Omega_{1})(w-\Omega_{1})}(v_{1}).$
@@
As for the relation (2.12), observe that due to the second part of Proposition 5.7 and 5.8, the product
@@
where $e^{\hbar}_{k+l}$ is the symmetrizer of Section 2.2. However, upon extending the scalars to $\mathbf{Z}_{\mathfrak{gl}_{N}}$, this is exactly the product of the shuffle algebra $\mathcal{F}^{\hbar}$ of 2.8, in other words, the “quotient” map
@@
is an algebra map. By Theorem 2.18, there is an algebra isomorphism $Y(\mathfrak{gl}_{2})^{+}\xrightarrow{\sim}\mathcal{F}^{\hbar}$. Twisting the resulting composition $Y(\mathfrak{gl}_{2})^{+}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ by negative sign (which does not influence the defining relations) sends $x_{+}(u)$ to $-\frac{1}{u-\Omega_{1}}(v_{1})$. ∎

### v2

@@ 4. Parabolic restriction and difference operators
In this section, we recall the notion of the parabolic restriction for Harish-Chandra bimodules essentially following [KS22], as well as the classical construction of Mickelsson algebras [Zhe90]. In particular, we identify an example of the latter with a negative half of the algebra of difference operators.
@@ 4.2. Parabolic restriction
In this subsection, we recall the notion of parabolic restriction as in [KS22] and explicitly compute in some concrete examples.
@@
In fact, one can explicitly compute the parabolic restriction as follows. By [AST71], there exists an element $P$ of $\mathfrak{h}_{N}$-weight zero in a certain completion of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}[\hbar^{-1}]$ whose action is well-defined on any right $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}$-module $F$ with a locally-finite action of $\mathfrak{n}_{-}$ satisfying $F\mathfrak{n}_{-}\subset\hbar F$; this element is called the extremal projector. It satisfies the following properties:
(4.6) $P\in 1+\mathfrak{n}_{-}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\cap\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\mathfrak{n}_{+},\qquad E_{ij}P=PE_{ji}=0,\ 1\leq i<j\leq N,\qquad P^{2}=P.$
@@
Usually, the cited results use the positive version of the parabolic restriction $X\mapsto(X^{\mathrm{gen}}/\mathfrak{n}_{+})^{N_{+}}$. For Miura transform of Section˜6, the negative one will be more convenient as the former is a map of left $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-modules; however, both versions are equivalent thanks to the property $\varpi_{N}(P)=P$. For instance, assume $\varpi_{N}$ can be extended to an isomorphism of a Harish-Chandra bimodule $X$. Then there is an anti-isomorphism
@@
[…]
@@
This is the standard Harish-Chandra isomorphism $\mathbf{Z}_{\mathfrak{gl}_{N}}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{W}$, see [Hum08, Theorem 1.10].
@@ 4.3. Vector representation
where $\{j_{1}\ldots j_{r}\}$ is the complementary subset to $\{i_{1},\ldots,i_{s}\}$ inside $\{i+1,\ldots,N\}$. According to formulas [Mol02, (3.2),(3.3)], the trick of Example˜3.6, and Remark˜4.3, those vectors are indeed invariant under $N_{-}$ and satisfy
@@
Using the parabolic restriction, one can deduce the following explicit description of the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$ as bimodule over the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$. Before the statement, we need the following lemma which is essentially a part of [KNV11, Theorem 2]; for reader’s convenience, we repeat its proof in our notations.
@@
[…]
@@
The equality in the following proposition is understood as in Remark˜4.8.
@@
Observe that both sides are actions of certain elements of $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})$. By ˜4.5, it is enough to show the equality under the parabolic restriction functor. It follows from ˜4.6 that
@@
At the same time, recall that by Example˜4.4, we have $PA(u)=\prod_{j=1}^{N}(u-w_{j})$, therefore,
@@
Strictly speaking, we understand the equality of Proposition˜4.7 in the sense of power series expansion in $u^{-1}$. The right-hand side can be presented as
@@
while the left-hand side can be equivalently presented as conjugation by $\mathrm{qdet}(T(-u+N\hbar-\hbar))$ in the notations of Section˜1.2 which also defines an element of $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})[[u^{-1}]]$. Another way is via the coefficients (1.11): for any $v\in\mathbf{C}^{N}$, we have
@@
In fact, using ˜4.6 and (4.9), one can re-deduce the Cayley-Hamilton theorem of Proposition˜3.7: observe that for every $i$, we have
@@ 4.4. Mickelsson algebras and difference operators
By [KS22, Corollary 4.18], the natural lax monoidal structure on $\mathrm{res}$ given by tensor multiplication
@@
[Zhe90] Let $A\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$ be an algebra object. The Mickelsson algebra associated to $A$ is the $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodule $A^{\mathrm{gen}}/\!\!/N_{-}$ with the multiplication given by
@@
where the target is the space of polynomials of total degree $k$ in $\mathbf{u}^{-1}$ variables. By ˜3.9, the source has an action of the spherical subalgebra $e^{\hbar}_{k}\mathfrak{H}^{\hbar}_{k}e^{\hbar}_{k}$ of the degenerate affine Hecke algebra.
@@
Recall that $\mathbf{u}_{i_{1}}^{-1}\ldots\mathbf{u}_{i_{k}}^{-1}$ is proportional to $Pv_{i_{1}}P\ldots Pv_{i_{k}}P$. By ˜3.9, the action of a monomial on the latter is
@@
(it is well-defined once the function is symmetric). The result follows by ˜4.6. ∎
@@ 5. Kostant-Whittaker reduction and Yangians
In this section, we will study a variant of the constructions of Section˜4. Our main goal is to relate the shifted Yangian of $\mathfrak{sl}_{2}$ to the Whittaker analog of the Mickelsson algebra for $S^{\bullet}\mathbf{C}^{N}$.
@@
For any $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$, denote by $X/\mathfrak{n}_{+}^{\psi}$ the quotient by the right action of $\mathfrak{n}_{+}^{\psi}$. Recall that by Kostant’s construction [Kos78], the space of invariants $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\!\!/_{\psi}\,N_{+}:=(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})/\mathfrak{n}_{+}^{\psi})^{N_{+}}$ is isomorphic to the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$. It admits the following categorification.
@@
[BF08] The Kostant-Whittaker reduction is a functor
@@
There is a Whittaker analog of the extremal projector (4.6) introduced in [Kal25] under the name of the Kirillov projector.
@@
$\begin{split}(x-\psi(x))P_{\psi}&=0,\ x\in\mathfrak{n}_{+},\\ P_{\psi}(E_{ij}+\delta_{ij}\hbar(N-i))&=0,\ 1\leq j\leq i\leq N-1,\\ (P_{\psi})^{2}=P_{\psi},&\qquad P_{\psi}\Big{|}_{(X/\mathfrak{n}_{+}^{\psi})^{N_{+}}}=\mathrm{id}.\end{split}$
@@
In the notations of [Kal25, Definition 5.2], we can apply (1.6) to $P_{\mathfrak{m}_{N}}^{\psi}(N\hbar-\hbar,\ldots,\hbar)$; by [Kal25, Theorem 5.3], it satisfies the properties of the theorem for $\psi(E_{ij})=\delta_{i+1,j}$. For general $\psi\in\mathbf{C}^{*}$, we can apply the automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$ defined by $x\mapsto x/\psi$ for $x\in\mathfrak{gl}_{N}$ and $\hbar\mapsto\hbar/\psi$. ∎
As in Section˜4, we use the following.
@@
We can compute explicitly this trivialization for the vector representation $\mathbf{C}^{N}$ using the canonical homomorphism (3.3). The following can be proved analogously to [Kal25, Proposition 6.18].
@@
Combining (5.3), ˜5.4 and Proposition˜3.7, we conclude that the map
@@
is an isomorphism of modules over $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]$. By Proposition˜4.7, we have
@@
By expanding in powers of $u$ as in Remark˜4.8, we can conclude that an endomorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ commuting with the left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-action commutes with the right $\mathbf{Z}_{\mathfrak{gl}_{N}}$-action if and only if it commutes with $\Omega$. Therefore, such an endomorphism commutes with the algebra action of $\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))$, giving a canonical isomorphism
@@ 5.1. Kostant-Mickelsson-Whittaker algebras
see [BF08]. In particular, we have an analog of ˜4.10.
@@
Recall the action of the degenerate affine Hecke algebra $\mathfrak{H}^{\hbar}_{k}$ on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$ from Section˜3.1. It descends to an action on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}/\!\!/_{\psi}\,N_{+}$. Recall the central elements $A(u)=\sum_{i}A_{i}u^{N-i}$ of (1.11). Define the elements $A^{(j)}_{i}\in\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega_{1},\ldots,\Omega_{k}]$ for $1\leq j\leq k+1$ inductively via $A^{(1)}_{i}=A_{i}$ for all $i$ and
@@
Using Remark˜4.8, it can be rewritten as
@@
The following is a generalization of Proposition˜5.5.
@@
By Proposition˜4.7, the right action of the center on each tensor component satisfies
@@
By ˜A.2, they generate the ring of symmetric functions in $\{\Omega_{i}\}$. Therefore, we have a desired isomorphism. ∎
@@
The following theorem relates the positive part of the shifted Yangian from ˜2.13 to the Kostant-Mickelsson-Whittaker algebra associated to $S^{\bullet}\mathbf{C}^{N}$. Observe similarity of the explicit formulas below to the quantization of Zastava spaces for $\mathrm{SL}_{2}$ constructed in [FR14].
@@
$A(z)\frac{1}{w-\Omega_{1}}(v_{1})+\frac{1}{w-\Omega_{1}}(v_{1})A(z)=-\frac{\hbar}{z-w}A(z)\left(\frac{1}{z-\Omega_{1}}-\frac{1}{w-\Omega_{1}}\right)(v_{1}).$
@@
and, due to Proposition˜4.7, the left-hand side is
$A(z)\frac{1}{w-\Omega_{1}}(v_{1})+A(z)\frac{1}{w-\Omega_{1}}(v_{1})\left(1+\frac{\hbar}{z-\Omega_{1}}\right)=A(z)\frac{\hbar}{(z-\Omega_{1})(w-\Omega_{1})}(v_{1}).$
@@
As for the relation (2.12), observe that due to the second part of Proposition˜5.7 and ˜5.8, the product
@@
where $e^{\hbar}_{k+l}$ is the symmetrizer of Section˜2.2. However, upon extending the scalars to $\mathbf{Z}_{\mathfrak{gl}_{N}}$, this is exactly the product of the shuffle algebra $\mathcal{F}^{\hbar}$ of ˜2.8, in other words, the “quotient” map
@@
is an algebra map. By Theorem˜2.18, there is an algebra isomorphism $Y(\mathfrak{gl}_{2})^{+}\xrightarrow{\sim}\mathcal{F}^{\hbar}$. Twisting the resulting composition $Y(\mathfrak{gl}_{2})^{+}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}$ by negative sign (which does not influence the defining relations) sends $x_{+}(u)$ to $\frac{1}{u-\Omega_{1}}(v_{1})$. ∎